OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 3x2 – 48x + 153 = 0

OpenStudy (paxpolaris):

divide by 3:\[x^2-16x+51=0\]

OpenStudy (paxpolaris):

then move 51 to the other side: \[x^2-16x=-51\]

OpenStudy (paxpolaris):

then add a number to both sides so that the Left Hand Side is a perfect square: \[x^2-16x+C= -51+C\]

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (paxpolaris):

\[C=\left( -16 \over 2 \right)^2\]

OpenStudy (paxpolaris):

\[x^2-16x+64=-51+64\]\[\left( x-8 \right)^2=13\]\[x-8=\pm \sqrt{13}\]\[x=8\pm \sqrt{13}\]